OpenStudy (anonymous):

what is the x-coordinate of the solution of system of equations? 2x-y=6 y=-3x+19

OpenStudy (anonymous):

rearrange the equation of 2x-y=6 into terms of y first.

OpenStudy (anonymous):

so y=-2x+6?

OpenStudy (anonymous):

y=2x+6

OpenStudy (anonymous):

2x-y=6 (bring the y to the right hand side so + y on both sides) and also (bring the six to the left hand side by -6 on both sides) 2x-6=y y=2x-6

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now, solve both equations

OpenStudy (anonymous):

1) y=2x-6 2) y=-3x+19

OpenStudy (anonymous):

tell me what you get for the x value afterwards.

OpenStudy (anonymous):

how though?

OpenStudy (anonymous):

do you know how to solve for both equations?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

okok, let me explain sonja, it's sonja right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1) y=2x-6 2) y=-3x+19 we want to find the x value and to do this we have to get rid of the y value. Since we rearranged the first equation already in terms of y, we can directly solve for x now.

OpenStudy (anonymous):